# Three meta-questions on infinite-domain Constraint Satisfaction Problems

### v1

@@ 5.2. Our sandwich recipe
Let $\mathfrak{A}$ and $\mathfrak{B}$ be two countable homogeneous structures without algebraicity over disjoint finite relational signatures $\tau$ and $\sigma$. Then there exists an up to isomorphism unique homogeneous $(\tau\cup\sigma)$-structure $\mathfrak{A}\ast\mathfrak{B}$, called their generic superposition, with the property that $\mathfrak{A}\ast\mathfrak{B}|_{\tau}$ and $\mathfrak{A}\ast\mathfrak{B}|_{\sigma}$ are isomorphic to $\mathfrak{A}$ and $\mathfrak{B}$, respectively. If both $\mathfrak{A}$ and $\mathfrak{B}$ are finitely bounded, then $\mathfrak{A}\ast\mathfrak{B}$ is finitely bounded as well.
Proof of Theorem 2.6.
Let $\mathfrak{A}$ be a non-trivial reduct of a finitely bounded homogeneous structure $\mathfrak{B}$. In the first step, we take the structures $\mathbb{Q}^{\mathfrak{A}}$ and $(\mathbb{Q};<)^{\mathfrak{B}}$ from Theorem 2.2 to remove algebraicity from $\mathfrak{A}$ and $\mathfrak{B}$ while ensuring that $\mathfrak{\mathbb{Q}^{\mathfrak{A}}}$ pp-constructs $\mathfrak{K}_{3}$ if and only if $\mathfrak{A}$ does and $\mathrm{CSP}(\mathfrak{A})$ and $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$ are Datalog-interreducible. Next, we take the generic superposition of $(\mathbb{Q};<)^{\mathfrak{B}}$ with $(\mathbb{Q};<)$, which exists by Proposition 5.3. We set $\widehat{\mathfrak{B}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(\mathbb{Q};<)^{\mathfrak{B}}\ast(\mathbb{Q};<)$ and define $\widehat{\mathfrak{A}}$ as the reduct of $\widehat{\mathfrak{B}}$ to the signature of $\mathbb{Q}^{\mathfrak{A}}$. By Proposition 5.3, $\widehat{\mathfrak{B}}$ is finitely bounded homogeneous and $\widehat{\mathfrak{A}}$ and $\mathbb{Q}^{\mathfrak{A}}$ are isomorphic. But now $\widehat{\mathfrak{A}}$ and $\widehat{\mathfrak{B}}$ satisfy the prerequisites of Proposition 5.2. It follows that $\mathfrak{A}^{\prime}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}={\widehat{\mathfrak{A}}}^{[d]}$, $\mathfrak{S}_{1}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}={\mathfrak{S}}^{[d]}$, and $\mathfrak{S}_{2}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}={\widehat{\mathfrak{A}}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\widehat{\mathfrak{B}})}}}$ for $\mathfrak{S}$ as in Proposition 5.2 witness items 14 of the theorem. Item 5 follows directly from Theorem 2.2(1) and Proposition 5.1(5). Similarly, item 6 follows directly from Theorem 2.2(2), Proposition 5.1(7), and [15, Theorem 1.5]. ∎
6. Conclusion
6.1. Necessity of preprocessing
In our proof of Theorem 2.6, we apply Proposition 5.2 to a structure which we previously “preprocessed” using Theorem 2.2 (removing algebraicity) and generic superpositions. It is natural to ask how strong Proposition 5.2 is on its own: can a result similar to Theorem 2.6 be proved directly using this proposition, perhaps under a more general structural assumption, e.g. that $\mathfrak{A}$ is a model-complete core? Here the answer seems to be negative. Consider the model-complete core structure $\mathfrak{A}=(\mathbb{Q};<,\{(x,y,z)\in\mathbb{Q}^{3}\mid x\geq y\text{ or }x\geq z\})$ from the scope of Conjecture 1.4 ($\mathfrak{A}_{2}$ in Example 4.6); its CSP is polynomial-time tractable [23]. Let $\mathfrak{S}$ be any finite structure for which there exists a homomorphism $h\colon\mathfrak{S}\rightarrow{\mathfrak{A}}^{[d]}$, and let $\mathfrak{B}$ be an arbitrary $\omega$-categorical expansion of $\mathfrak{A}$. Let $A_{S}$ be the set of the elements in $A$ that appear as an entry of some element of $h(S)$, and let $\mathfrak{A}_{S}$ be the substructure of $\mathfrak{A}$ on $A_{S}$. Since ${\mathfrak{A}}^{[d]}\to\smash{{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}$, the structure ${\mathfrak{A}}^{[d]}_{S}$ is a finite sandwich for $\mathrm{PCSP}(\mathfrak{S},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$. We show that $\mathrm{CSP}(\smash{{\mathfrak{A}}^{[d]}_{S}})$ is tractable, and hence this PCSP is finitely tractable (unless P=NP, in which case this entire discussion is pointless). To this end, observe that the binary minimum operation $\min(x,y)$ is a polymorphism of $\mathfrak{A}$. This operation is conservative: $\min(x,y)\subseteq\{x,y\}$ holds for all $x,y$. Hence, its restriction to $A_{S}^{2}$ induces a polymorphism of $\mathfrak{A}_{S}$, which itself induces a polymorphism of ${\mathfrak{A}}^{[d]}_{S}$ through its component-wise action. Since this operation is cyclic, ${\mathfrak{A}}^{[d]}_{S}$ does not pp-construct $\mathfrak{K}_{3}$ [9, Theorem 1.4], and hence we are done by Theorem 1.1. This issue cannot be fixed simply by taking an expansion of $\mathfrak{A}$ by the inequality relation $\neq$ because $\mathrm{CSP}(\mathfrak{A};\neq)$ is NP-complete [23].
6.2. Numeric algorithms fail
An interesting open question is whether the PCSP templates generated by Theorem 2.6 can be solved by some universal algorithm. Many known PCSP problems admitting infinite tractable sandwiches can be solved by numeric relaxation algorithms such as BLP (the basic linear programming relaxation), AIP (the basic affine integer relaxation) or BLP+AIP (a combination of both) (see, e.g., [6, 35]). The proof of Proposition 5.2 shows that none of the PCSP templates produced by Theorem 2.6 admits a cyclic polymorphism. Since solvability of PCSPs by BLP is characterized by admitting totally symmetric polymorphisms of all arities [6, Theorem 7.9], which would also be cyclic, BLP does not solve any of those PCSPs.
Regarding the combination BLP+AIP, we remark that the construction of $\mathbb{Q}^{\mathfrak{A}}$ and $(\mathbb{Q};<)^{\mathfrak{B}}$ in the proof of Theorem 2.2 can be modified so that, for some finite substructure $\mathfrak{S}$ of $\mathbb{Q}^{\mathfrak{A}}$, there is no cyclic and no 2-block symmetric operation in $\mathrm{Pol}(\mathfrak{S},\mathbb{Q}^{\mathfrak{A}})$ (all other properties of $\mathbb{Q}^{\mathfrak{A}}$ and $(\mathbb{Q};<)^{\mathfrak{B}}$ remain preserved). Similarly to cyclic identities, also any 2-block symmetric identity can be lifted from $\mathrm{Pol}({\mathfrak{S}}^{[d]},{(\mathbb{Q}_{\!{}_{\vphantom{/}\smash{/G}}}^{\mathfrak{A}})}^{[d]})$ to $\mathrm{Pol}(\mathfrak{S},\mathbb{Q}^{\mathfrak{A}})$, where $G\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\smash{\mathrm{Aut}(\mathbb{Q}^{\mathfrak{B}}\ast(\mathbb{Q};<))}$. Thus, since solvability of a finite-domain PCSP by BLP+AIP is characterized by the existence of 2-block symmetric polymorphisms [35, Theorem 4], $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{(\mathbb{Q}_{\!{}_{\vphantom{/}\smash{/G}}}^{\mathfrak{A}})}^{[d]})$ is not solvable by BLP+AIP. For details, we refer the reader to [59, Theorem 2], where this idea was used in concrete cases of non-finitely tractable PCSPs. It follows that we can exclude BLP and BLP+AIP as potential algorithms solving the PCSPs produced by Theorem 2.6.
6.3. Topology is irrelevant
@@ A.1. Omitted definitions
We expand on some of the definitions appearing in the paper that were, for the sake of brevity, not presented in much detail. We also add the concept of a relational congruence, which will be useful in the the presentation of the proofs of Proposition 4.5, Lemma B.4, and Theorem 2.2, and state the theorem of Kechris, Pestov, and Todorčević referred to in Section 2.4, which is needed in the proof of Proposition 4.5.
Let $\mathfrak{A}$ be a structure over a finite relational signature containing $\tau$. An equivalence relation $E$ over $A$ is a relational congruence on $\mathfrak{A}$ with respect to $\tau$ if the definitions of the corresponding relations of $\mathfrak{A}/E$ do not depend on the choices of the representatives of the equivalence classes of $E$: for every $R\in\tau$ and all tuples $\bar{t}$ we have $q_{E}(\bar{t})\in R^{\mathfrak{A}/{E}}$ if and only if $\bar{t}\in R^{\mathfrak{A}}$.
Recall from the introduction that an $\omega$-categorical structure $\mathfrak{A}$ is a model-complete core if the endomorphisms of $\mathfrak{A}$ preserve all orbits of $\mathrm{Aut}(\mathfrak{A})$. A reformulation of this definition that will be more convenient for us is that for every $e\in\mathrm{End}(\mathfrak{A})$ and every finite $F\subseteq A$, there exists $\alpha\in\mathrm{Aut}(\mathfrak{A})$ such that $e|_{F}=\alpha|_{F}$ [16]; we will use this in the proof of Theorem 2.2.
Next, we properly define the Ramsey property. For structures $\mathfrak{A}$ and $\mathfrak{W}$, we denote by $\binom{\mathfrak{W}}{\mathfrak{A}}$ the set of all embeddings of $\mathfrak{A}$ into $\mathfrak{W}$. A class $\mathcal{K}$ of structures over a common signature $\tau$ has the Ramsey property (RP) if, for all $\mathfrak{X},\mathfrak{Y}\in\mathcal{K}$ and $k\in\mathbb{N}$, there exists $\mathfrak{W}\in\mathcal{K}$ such that, for every map $f\colon\smash{\binom{\mathfrak{W}}{\mathfrak{X}}}\rightarrow[k]$, there exists $e\in\smash{\binom{\mathfrak{W}}{\mathfrak{Y}}}$ such that $f$ is constant on the set $\{e\circ u\;|\;u\in\textstyle\smash{\binom{\mathfrak{Y}}{\mathfrak{X}}}\}\subseteq\textstyle\smash{\binom{\mathfrak{W}}{\mathfrak{X}}}.$ It is a folklore fact that considering the special case $k=2$ is enough for verifying the RP [51], and we call any $\mathfrak{W}$ satisfying the above in this case a Ramsey witness for $(\mathfrak{X},\mathfrak{Y})$.
A homogeneous structure is Ramsey if its age has the Ramsey property. The age of a structure $\mathfrak{B}$, denoted by $\mathrm{age}(\mathfrak{B})$, is the class of all finite structures which embed into $\mathfrak{B}$.
The KPT correspondence.
@@
Let $\mathfrak{B}$ be a CSP-injective $\omega$-categorical model-complete core. We show that $\mathfrak{B}$ has no algebraicity. Suppose, on the contrary, that there exists $\bar{b}\in B^{k}$ and $b^{\prime}\notin\bar{b}$ such that the unary relation $\{b^{\prime}\}$ is first-order definable in $\mathfrak{B}$ using $\bar{b}$ as parameters. By [16, Theorem 4.5.1 and Lemma 4.5.5], the expansion $(\mathfrak{B},\{\bar{b}\})$ is an $\omega$-categorical model-complete core, and there exists a primitive positive formula $\phi(x)$ defining $\{b^{\prime}\}$ in $(\mathfrak{B},\{\bar{b}\})$. Also, by [16, Theorem 4.5.1], there exists a primitive positive formula $\psi(\bar{y})$ defining the orbit of $\bar{b}$ in $\mathfrak{B}$. We may assume that $\psi$ does not contain any equality atoms, otherwise we eliminate them by identifying variables. Let $\phi^{\prime}$ be the primitive positive formula obtained by replacing the constants $\bar{b}$ in $\phi$ with the variables $\bar{y}$, and set $\phi^{\prime\prime}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\psi(\bar{y})\wedge\phi^{\prime}(x_{1})\wedge\phi^{\prime}(x_{2})$; we may again assume that $\phi^{\prime}$ and $\phi^{\prime\prime}$ do not contain any equality atoms. We can use $\phi^{\prime\prime}$ to create an instance $\mathfrak{A}$ of $\mathrm{CSP}(\mathfrak{B})$ that maps homomorphically to $\mathfrak{B}$ but has no injective such homomorphism. More specifically, the domain of $\mathfrak{A}$ consists of the variables of $\phi^{\prime\prime}$, and the relations are determined by the atomic subformulas of $\phi^{\prime\prime}$. Clearly, $\mathfrak{A}$ has a homomorphism to $\mathfrak{B}$, because we can substitute $\bar{b}$ for $\bar{y}$ in $\phi^{\prime\prime}$ and $b$ for both $x_{1}$ and $x_{2}$. However, $\mathfrak{A}$ has no injective homomorphism to $\mathfrak{B}$. Indeed, if this was the case, then there would exist distinct witnesses $\bar{y}$, $x_{1}$, and $x_{2}$ for the satisfaction of $\phi^{\prime\prime}$ in $\mathfrak{B}$. By definition, there would exist $\alpha\in\mathrm{Aut}(\mathfrak{B})$ with $\alpha(\bar{y})=\bar{b}$. But then $\alpha(x_{1})=\alpha(x_{2})=c$, a contradiction to the injectivity of $\alpha$. We conclude that $\mathfrak{A}$ does not have any injective homomorphism to $\mathfrak{B}$, and hence $\mathfrak{B}$ is not CSP-injective, a contradiction to our original assumption. Hence, $\mathfrak{B}$ has no algebraicity. ∎
A.3. A proof of Proposition 4.2
@@
Let $\mathfrak{B}$ be an $\omega$-categorical Pol-injective structure. Note that every essentially injective operation preserves the relation $I_{4}$. Hence, every polymorphism of $\mathfrak{B}$ preserves $I_{4}$, and so $I_{4}$ is pp-definable in $\mathfrak{B}$.
Let $\phi(x,y,u,v)$ be a primitive positive formula that defines $I_{4}$ in $\mathfrak{B}$. Since $I_{4}$ contains all $4$-tuples with pairwise distinct entries, we may assume that $\phi$ contains no equality atoms, otherwise we eliminate them by identifying variables. But now we can use $\phi$ to create an instance $\mathfrak{A}$ of $\mathrm{CSP}(\mathfrak{B})$ that maps homomorphically to $\mathfrak{B}$ but has no injective such homomorphism. More specifically, the domain of $\mathfrak{A}$ consists of the variables of $\phi$, except that we identify $x$ and $y$ (but not $u$ and $v$), and the relations are determined by the atomic subformulas of $\phi$. We conclude that $\mathfrak{B}$ is not CSP-injective. ∎
A.4. A proof of Proposition 4.3
@@
The Pol-injectivity of $\mathfrak{A}_{I_{4}}$ follows from the known fact that every polymorphism of $I_{4}$ is essentially injective [17] (see also [16, Lemma 7.5.1]). We proceed with the proof of Datalog-interreducibility.
@@
Now suppose that there exists a homomorphism $g\colon\mathfrak{J}\rightarrow\mathfrak{A}_{I_{4}}$. Note that, by definition, for every $k$-ary $R\in\tau$, we have $([s_{i_{1}}]_{\approx},\dots,[s_{i_{k}}]_{\approx})\in R^{\mathcal{I}^{\ast}(\mathfrak{J})}$ if and only if there are $j_{1}\in[s_{i_{1}}]_{\approx},\dots,j_{k}\in[s_{i_{k}}]_{\approx}$ such that $(j_{1},\dots,j_{k})\in R^{\mathfrak{J}}$. By the definition of $\approx$, every function preserving $I_{4}$ must identify $\approx$-equivalent elements, and therefore $(g(s_{i_{1}}),\dots,g(s_{i_{k}}))=(g(j_{1}),\dots,g(j_{k}))$, which implies $(g(s_{i_{1}}),\dots,g(s_{i_{k}}))\in R^{\mathfrak{A}_{I_{4}}}$. Therefore, $g$ is a homomorphism from $\mathcal{I}^{\ast}(\mathfrak{J})$ to $\mathfrak{A}$.
@@
For (2), we show that fixing any tuple $\bar{t}$ of parameters does not yield any nontrivial fixed point of the automorphism group when stabilizing $\bar{t}$. Let $(a^{\prime},b^{\prime})$ be any element outside $\bar{t}$. Let $A^{\prime}$ be the set of all $a\in A$ such that $(a,b)$ appears in $\bar{t}$ for some $b\in B$. Move $a^{\prime}$ to a different element via some $g\in\mathrm{Aut}(\mathfrak{A})$ whilst fixing $A^{\prime}$; this is possible because $\mathfrak{A}$ has no algebraicity. Then $\left((g_{b})_{b\in B},id\right)$ with $g_{b}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=g$ if $b=b^{\prime}$ and $id$ otherwise is an automorphism of $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})$ fixing each entry in $\bar{t}$ and moving $(a^{\prime},b^{\prime})$.

### v2

@@ 5.2. Our sandwich recipe
[…]
Proof sketch (Theorem 2.6).
[…]
We select a finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ with $|S|\geq 3$ large enough so that every finite structure with a homomorphism to a substructure of $\mathfrak{A}$ of size $\leq n$ also has a homomorphism to $\mathfrak{S}$; let $\widehat{\mathfrak{S}}$ be the substructure of $\widehat{\mathfrak{A}}$ induced by $\{0\}\times S$. We claim that $\mathfrak{A}^{\prime}\coloneqq{\widehat{\mathfrak{A}}}^{[d]}$, $\mathfrak{B}^{\prime}\coloneqq{\widehat{\mathfrak{B}}}^{[d]}$, $\mathfrak{S}_{1}\coloneqq{\widehat{\mathfrak{S}}}^{[d]}$, and $\mathfrak{S}_{2}\coloneqq{\widehat{\mathfrak{A}}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\widehat{\mathfrak{B}})}}}$ for $d$ as in Proposition 5.2 witness items 17 of the theorem. Items 1 to 6 follow from Theorem 2.2, Proposition 5.1, and Proposition 5.2. For item 7, note that if $\mathfrak{X}$ maps homomorphically to a substructure of $\mathfrak{A}$ of size $\leq n$, then $\mathfrak{X}\rightarrow\mathfrak{S}$, hence $\mathcal{I}(\mathfrak{X})\rightarrow\widehat{\mathfrak{S}}$ and hence $\mathcal{J}(\mathfrak{X})={\mathcal{I}(\mathfrak{X})}^{[d]}\rightarrow{\widehat{\mathfrak{S}}}^{[d]}=\mathfrak{S}_{1}$. Conversely, if $\mathfrak{X}\nrightarrow\mathfrak{A}$, then $\mathcal{I}(\mathfrak{X})\nrightarrow\widehat{\mathfrak{A}}$ by Theorem 2.2. One can show that if ${\mathcal{I}(\mathfrak{X})}^{[d]}\rightarrow\mathfrak{S}_{2}$, then also $\mathcal{I}(\mathfrak{X})\rightarrow\widehat{\mathfrak{A}}$, hence $\mathcal{J}(\mathfrak{X})\nrightarrow\mathfrak{S}_{2}$. ∎
6. Conclusion and outlook
6.1. Topology seems irrelevant
@@
6.2. Necessity of cheese preprocessing
In our proof of Theorem 2.6, we apply Proposition 5.2 to a structure which we previously “preprocessed” using Theorem 2.2 (removing algebraicity) and generic superpositions. It is natural to ask how strong Proposition 5.2 is on its own: can a result similar to Theorem 2.6 be proved directly using this proposition, perhaps under a more general structural assumption, e.g. that $\mathfrak{A}$ is a model-complete core? Here the answer seems to be negative. Consider the model-complete core structure
$\mathfrak{A}=(\mathbb{Q};<,\{(x,y,z)\in\mathbb{Q}^{3}\mid x\geq y\text{ or }x\geq z\})$
within the scope of Conjecture 1.4 ($\mathfrak{A}_{2}$ in Example 4.6); its CSP is polynomial-time tractable [22]. Let $\mathfrak{S}$ be any finite structure for which there exists a homomorphism $h\colon\mathfrak{S}\rightarrow{\mathfrak{A}}^{[d]}$, and let $\mathfrak{B}$ be an arbitrary $\omega$-categorical expansion of $\mathfrak{A}$. Let $A_{S}$ be the set of the elements in $A$ that appear as an entry of some element of $h(S)$, and let $\mathfrak{A}_{S}$ be the substructure of $\mathfrak{A}$ on $A_{S}$. Since $\smash{{\mathfrak{A}}^{[d]}\to{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}$, the structure $\smash{{\mathfrak{A}}^{[d]}_{S}}$ is a finite cheese for $\smash{\mathrm{PCSP}(\mathfrak{S},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})}$.
We claim that $\mathrm{CSP}({\mathfrak{A}}^{[d]}_{S})$ is tractable, and hence this PCSP is finitely tractable. To see this, observe that the binary minimum operation $\min(x,y)$ is a polymorphism of $\mathfrak{A}$. This operation is conservative: $\min(x,y)\subseteq\{x,y\}$ holds for all $x,y$. Hence, its restriction to $A_{S}^{2}$ induces a polymorphism of $\mathfrak{A}_{S}$, which itself induces a polymorphism of $\smash{{\mathfrak{A}}^{[d]}_{S}}$ through its component-wise action. Since this operation is cyclic, $\smash{{\mathfrak{A}}^{[d]}_{S}}$ does not pp-construct $\mathfrak{K}_{3}$ [9, Theorem 1.4], and hence we are done by Theorem 1.1. This issue cannot be fixed simply by taking an expansion of $\mathfrak{A}$ by the inequality relation $\neq$ because $\mathrm{CSP}(\mathfrak{A};\neq)$ is NP-complete [22].
6.3. Algorithmic properties of our sandwiches
An interesting open question is whether the PCSP templates generated by Theorem 2.6 can be solved by some universal algorithm. Many known PCSP problems admitting infinite tractable cheeses can be solved by numeric relaxation algorithms such as BLP (the basic linear programming relaxation), AIP (the basic affine integer relaxation) or BLP+AIP (a combination of both) (see, e.g., [6, 34]). The proof of Proposition 5.2 shows that none of the PCSP templates produced by Theorem 2.6 admits a cyclic polymorphism. Since solvability of PCSPs by BLP is characterized by admitting totally symmetric polymorphisms of all arities [6, Theorem 7.9], which would also be cyclic, BLP does not solve any of those PCSPs. Regarding the combination BLP+AIP, we remark that replacing $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $\smash{\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}}$ in the proof of Theorem 2.6 by $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ and $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}$ ensures that there is no cyclic and no 2-block symmetric operation in $\mathrm{Pol}(\mathfrak{S},(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$. In fact, we get the following much stronger statement, which follows directly from the fact that we take an expansion by $I_{4}$.
Proposition 6.1 (cf. proof of Lemma 7.5.1 in [15]).
Let $\Sigma$ be any finite height-1 condition that cannot be satisfied by any set of essentially injective operations from an at least 2-element set $S_{1}$ to a countably infinite set $S_{2}$. Then, for every at least 2-element substructure $\mathfrak{S}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$, we have that $\mathrm{Pol}(\mathfrak{S},(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ does not satisfy $\Sigma$. ∎
[…]
The lifting method we and Mottet [57] use provably cannot be extended to all height-1 conditions covered by Proposition 6.1. To see this, note that the Olšák identities
$f(y,x,x,x,y,y)\approx f(x,y,x,y,x,y)\approx f(x,x,y,y,y,x)$
cannot be satisfied by essentially injective operations but every finite-domain PCSP template without an Olšák polymorphism must be NP-hard [6, Corollary 6.3]. This means that the Olšák identities cannot possibly be prevented in our PCSPs without losing solvability in polynomial time (unless P=NP).
@@ A.1. Omitted definitions
We expand on some of the definitions appearing in the paper that were, for the sake of brevity, not presented in much detail. We also add the concept of a relational congruence, which will be useful in the the presentation of the proofs of Proposition 4.5, Lemma B.2, and Theorem 2.2, and state the theorem of Kechris, Pestov, and Todorčević referred to below Proposition 4.5, which is needed in the proof of Proposition 4.5.
Let $\mathfrak{A}$ be a structure over a finite relational signature containing $\tau$. The age of $\mathfrak{A}$, denoted by $\mathrm{age}(\mathfrak{A})$, is the class of all finite structures which embed into $\mathfrak{A}$. An equivalence relation $E$ over $A$ is a relational congruence on $\mathfrak{A}$ with respect to $\tau$ if the definitions of the corresponding relations of $\mathfrak{A}/E$ do not depend on the choices of the representatives of the equivalence classes of $E$: for every $R\in\tau$ and all tuples $\bar{t}$ we have $q_{E}(\bar{t})\in R^{\mathfrak{A}/{E}}$ if and only if $\bar{t}\in R^{\mathfrak{A}}$.
Next, recall from the introduction that an $\omega$-categorical structure $\mathfrak{A}$ is a model-complete core if the endomorphisms of $\mathfrak{A}$ preserve all orbits of $\mathrm{Aut}(\mathfrak{A})$. A reformulation of this definition that will be more convenient for us is that for every $e\in\mathrm{End}(\mathfrak{A})$ and every finite $F\subseteq A$, there exists $\alpha\in\mathrm{Aut}(\mathfrak{A})$ such that $e|_{F}=\alpha|_{F}$ [15]; we will use this in the proof of Theorem 2.2.
Now, we properly define the Ramsey property. For structures $\mathfrak{A}$ and $\mathfrak{W}$, we denote by $\binom{\mathfrak{W}}{\mathfrak{A}}$ the set of all embeddings of $\mathfrak{A}$ into $\mathfrak{W}$. A class $\mathcal{K}$ of structures over a common signature $\tau$ has the Ramsey property (RP) if, for all $\mathfrak{X},\mathfrak{Y}\in\mathcal{K}$ and $k\in\mathbb{N}$, there exists $\mathfrak{W}\in\mathcal{K}$ such that, for every map $f\colon\binom{\mathfrak{W}}{\mathfrak{X}}\rightarrow[k]$, there exists $e\in\binom{\mathfrak{W}}{\mathfrak{Y}}$ such that $f$ is constant on the set $\{e\circ u\;|\;u\in\textstyle\binom{\mathfrak{Y}}{\mathfrak{X}}\}\subseteq\textstyle\binom{\mathfrak{W}}{\mathfrak{X}}.$ It is a folklore fact that considering the special case $k=2$ is enough for verifying the RP [49], and we call any $\mathfrak{W}$ satisfying the above in this case a Ramsey witness for $(\mathfrak{X},\mathfrak{Y})$. A homogeneous structure is Ramsey if its age has the Ramsey property.
A.2. The KPT correspondence.
@@
Let $\mathfrak{B}$ be a CSP-injective $\omega$-categorical model-complete core. We show that $\mathfrak{B}$ has no algebraicity. Suppose, on the contrary, that there exists $\bar{b}\in B^{k}$ and $b^{\prime}\notin\bar{b}$ such that the unary relation $\{b^{\prime}\}$ is first-order definable in $\mathfrak{B}$ using $\bar{b}$ as parameters. By [15, Theorem 4.5.1 and Lemma 4.5.5], the expansion $(\mathfrak{B},\{\bar{b}\})$ is an $\omega$-categorical model-complete core, and there exists a primitive positive formula $\phi(x)$ defining $\{b^{\prime}\}$ in $(\mathfrak{B},\{\bar{b}\})$. Also, by [15, Theorem 4.5.1], there exists a primitive positive formula $\psi(\bar{y})$ defining the orbit of $\bar{b}$ in $\mathfrak{B}$. We may assume that $\psi$ does not contain any equality atoms, otherwise we eliminate them by identifying variables. Let $\phi^{\prime}$ be the primitive positive formula obtained by replacing the constants $\bar{b}$ in $\phi$ with the variables $\bar{y}$, and set $\phi^{\prime\prime}\coloneqq\psi(\bar{y})\wedge\phi^{\prime}(x_{1})\wedge\phi^{\prime}(x_{2})$; we may again assume that $\phi^{\prime}$ and $\phi^{\prime\prime}$ do not contain any equality atoms. We can use $\phi^{\prime\prime}$ to create an instance $\mathfrak{A}$ of $\mathrm{CSP}(\mathfrak{B})$ that maps homomorphically to $\mathfrak{B}$ but has no injective such homomorphism. More specifically, the domain of $\mathfrak{A}$ consists of the variables of $\phi^{\prime\prime}$, and the relations are determined by the atomic subformulas of $\phi^{\prime\prime}$. Clearly, $\mathfrak{A}$ has a homomorphism to $\mathfrak{B}$, because we can substitute $\bar{b}$ for $\bar{y}$ in $\phi^{\prime\prime}$ and $b$ for both $x_{1}$ and $x_{2}$. However, $\mathfrak{A}$ has no injective homomorphism to $\mathfrak{B}$. Indeed, if this was the case, then there would exist distinct witnesses $\bar{y}$, $x_{1}$, and $x_{2}$ for the satisfaction of $\phi^{\prime\prime}$ in $\mathfrak{B}$. By definition, there would exist $\alpha\in\mathrm{Aut}(\mathfrak{B})$ with $\alpha(\bar{y})=\bar{b}$. But then $\alpha(x_{1})=\alpha(x_{2})=c$, a contradiction to the injectivity of $\alpha$. We conclude that $\mathfrak{A}$ does not have any injective homomorphism to $\mathfrak{B}$, and hence $\mathfrak{B}$ is not CSP-injective, a contradiction to our original assumption. Hence, $\mathfrak{B}$ has no algebraicity. ∎
A.4. A proof of Proposition 4.2
@@
Let $\mathfrak{B}$ be an $\omega$-categorical Pol-injective structure. Note that every essentially injective operation preserves the relation $I_{4}$. Hence, every polymorphism of $\mathfrak{B}$ preserves $I_{4}$, and so $I_{4}$ is pp-definable in $\mathfrak{B}$. Let $\phi(x,y,u,v)$ be a primitive positive formula that defines $I_{4}$ in $\mathfrak{B}$. Since $I_{4}$ contains all $4$-tuples with pairwise distinct entries, we may assume that $\phi$ contains no equality atoms, otherwise we eliminate them by identifying variables. But now we can use $\phi$ to create an instance $\mathfrak{A}$ of $\mathrm{CSP}(\mathfrak{B})$ that maps homomorphically to $\mathfrak{B}$ but has no injective such homomorphism. More specifically, the domain of $\mathfrak{A}$ consists of the variables of $\phi$, except that we identify $x$ and $y$ (but not $u$ and $v$), and the relations are determined by the atomic subformulas of $\phi$. We conclude that $\mathfrak{B}$ is not CSP-injective. ∎
A.5. A proof of Proposition 4.3
@@
The Pol-injectivity of $\mathfrak{A}_{I_{4}}$ follows from the known fact that every polymorphism of $I_{4}$ is essentially injective [16] (see also [15, Lemma 7.5.1]). We proceed with the proof of Datalog-interreducibility.
@@
Now suppose that there exists a homomorphism $g\colon\mathfrak{J}\rightarrow\mathfrak{A}_{I_{4}}$. Note that, by definition, for every $k$-ary $R\in\tau$, we have $(s_{i_{1}},\dots,s_{i_{k}})\in R^{\mathcal{I}^{\ast}(\mathfrak{J})}$ if and only if there are $j_{1}\in[s_{i_{1}}]_{\approx},\dots,j_{k}\in[s_{i_{k}}]_{\approx}$ such that $(j_{1},\dots,j_{k})\in R^{\mathfrak{J}}$. By the definition of $\approx$, every function preserving $I_{4}$ must identify $\approx$-equivalent elements, and therefore $(g(s_{i_{1}}),\dots,g(s_{i_{k}}))=(g(j_{1}),\dots,g(j_{k}))$, which implies $(g(s_{i_{1}}),\dots,g(s_{i_{k}}))\in R^{\mathfrak{A}_{I_{4}}}$. Therefore, $g$ is a homomorphism from $\mathcal{I}^{\ast}(\mathfrak{J})$ to $\mathfrak{A}$.
@@
For (2), we must show that fixing any tuple $\bar{t}$ of parameters does not yield any nontrivial fixed point of the automorphism group when stabilizing $\bar{t}$. Let $(a^{\prime},b^{\prime})$ be any element outside $\bar{t}$. Let $A^{\prime}$ be the set of all elements $a\in A$ such that $(a,b)$ appears in $\bar{t}$ for some $b\in B$. Move $a^{\prime}$ to a different element via some $g\in\mathrm{Aut}(\mathfrak{A})$ whilst fixing $A^{\prime}$; this is possible because $\mathfrak{A}$ has no algebraicity. Then $\left((g_{b})_{b\in B},id\right)$ with $g_{b}\coloneqq g$ if $b=b^{\prime}$ and $id$ otherwise is an automorphism of $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})$ fixing every entry in $\bar{t}$ and moving $(a^{\prime},b^{\prime})$.